ques:- The latent heat of vaporization of water is 2240 J. If the work done in the process of vaporization of 1 gm is 168 J, then increase in internal energy is?

Grade:12

2 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

following first law of thermodynamics

dQ = dU + dW

for 1 gm Lv = 224 J

delU = delQ - delW = 224 - 168 = 56J

i hope it explains !! :)

Anonymous
5 years ago
Given
Latent heat of vapourisation=+2240J 
Work done=168J
Using 1st Law of thermodynamics:-
2240=dU+168
dU=2072 J